Convert ((1−i)/2) into trigonometric form.


1
Expert's answer
2022-06-06T02:45:19-0400

"\frac{1-i}{2}=\frac{1}{\sqrt{2}}\big(\frac{1}{\sqrt{2}}-\frac{1}{\sqrt{2}}i\big)=\frac{1}{\sqrt{2}}\big(\cos(-\pi/4)+i\sin(-\pi/4)\big)"